OpenStudy (anonymous):

A steel block is extends a spring 14cm when hung from it. When drawn across a surface using the same spring at a constant speed, the spring extends 3 cm. What is the coefficient of friction? Help? We get friction using the formula F = u * mg and the extension of the cylinder seems to imply Hooke’s law is relevant; F = k * l.

OpenStudy (anonymous):

hanging: mg = kx = k*.14 k=mg/.14 dragging: kx = mu*mg k*.03 = mu*mg (mg/.14)*.03 = mu*mg .03/.14 = mu